### v2

@@ 2 Results
In both experiments, uncalibrated AI consistently hindered human-AI collaboration, and participants were unable to effectively calibrate their trust towards uncalibrated AI. Increased transparency, i.e., trust calibration support, can enhance the perception of AI’s uncalibrated state by participants and reduce misuse caused by overconfident AI; however, it introduces a distrust towards AI’s prediction accuracy as well as an increase of disuse, thereby not significantly improving the outcome of human-AI collaboration. All statistical test result details, descriptive data, tests related to control variables, and demographic data are provided in the supplementary material.
Figure 3: Heat maps illustrating participants’ perception of AI’s confidence calibration. Each row in a given map corresponds to different groups from EXP1 and EXP2, categorized by the confidence calibration states of the AI they collaborated with. Each column represents the AI’s confidence calibration states perceived by participants within that group, which could be unconfident, well-calibrated, or overconfident. The red boxes highlight instances where the perceived AI’s confidence calibration state matched the AI’s confidence calibration state in that group, indicating participants who correctly perceived the AI’s confidence calibration state. (a) is the result of EXP1. (b) is the result of EXP2.
Figure 4: Results of ANOVA in EXP1 are depicted in violin and box plots, with the dependent variables being Participants’ Adoption Rate, Disuse Rate, Misuse Rate, Accuracy Increase, Trust towards AI’s Confidence Score, Trust towards AI’s Prediction Accuracy, and Trust towards AI’s Overall Capability. The $p$-values of the AI confidence calibration states’ main effects from the one-way ANOVA are indicated below each figure, and the significance levels from post-hoc analyses are denoted within the figures (ns: $p>0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$). F-values with degrees of freedom are also reported below the figure. Differences between groups from post-hoc analyses, excluding comparisons between EXP1-U and EXP1-O, are marked in the figure. Further information, such as the verification of homogeneity assumptions, analytical methods, and effect sizes, can be found in the supplementary materials.
@@ 2.1.1 Perceiving AI’s Overconfidence and Unconfidence is Challenging
The results of the treatments EXP1-U and EXP1-O, as depicted on Fig. 3 (a), indicate that, it is difficult for the majority of participants to perceive if the provided AI’s confidence level aligns with its actual correctness likelihood. In EXP1-U, 28.571% of participants correctly perceived the AI’s unconfidence. In EXP1-O, only 26.190% of participants correctly perceived the overconfidence. The majority of participants deemed the AI’s confidence calibration state appropriate when the AI was either overconfident or unconfident, suggesting an overtrust in AI’s confidence.
@@ 2.1.4 Misuse and Disuse Further Impair Human-AI Collaboration Outcomes
The results of the mediation analysis in Fig. 6 (a) and (b) further revealed the relationships between uncalibrated AI confidence, participants’ misuse/disuse, participants’ adoption of AI decisions, and the accuracy increase without trust calibration support. Utilizing data from EXP1-W and EXP1-O, we found that the misuse rate mediated the influence of AI’s overconfidence on participants’ adoption rate ($\beta_{i}$ (indirect effect) $=6.875$, $p=0.009$; $\beta_{d}$ (direct effect) $=5.737$, $p=0.228$) and final accuracy increase ($\beta_{i}$ $=-4.673$, $p<0.001$; $\beta_{d}$ = $-0.011$, $p=0.994$). With data from EXP1-U and EXP1-W, it was discovered that the disuse rate mediated the effect of AI’s unconfidence on participants’ adoption rate ($\beta_{i}=-16.869$, $p<0.001$; $\beta_{d}=0.387$, $p=0.914$) and final accuracy increase ($\beta_{i}=-5.809$, $p<0.001$; $\beta_{d}=0.412$, $p=0.728$).
Figure 5: Results of ANOVA in EXP2 are depicted in violin and box plots, with the dependent variables being participants’ Adoption Rate, Disuse Rate, Misuse Rate, Accuracy Increase, Trust towards AI’s Confidence Score, Trust towards AI’s Prediction Accuracy, and Trust towards AI’s Overall Capability. The $p$-values of the AI confidence calibration states’ main effects from the one-way ANOVA are indicated below each figure, and the significance levels from post-hoc analyses are denoted within the figures (ns: $p>0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$). F-values with degrees of freedom are also reported below the figure. Differences between groups from post-hoc analyses, excluding comparisons between EXP2-U and EXP2-O, are marked in the figure. Further information, such as the verification of homogeneity assumptions, analytical methods, and effect sizes, can be found in the supplementary materials.
Figure 6: The diagrams above illustrates the results of mediation analysis (ns: $p>0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$). See more details in supplementary materials. (a) Mediation analysis on EXP1-W and EXP1-O about the effects of uncalibrated AI confidence on accuracy increase, mediated by adoption rate and misuse. (b) Mediation analysis on EXP1-U and EXP1-W about the effects of uncalibrated AI confidence on accuracy increase, mediated by adoption rate and disuse. (c) Mediation analysis on EXP2-W and EXP2-O about the effects of uncalibrated AI confidence on accuracy increase, mediated by adoption rate and disuse. (d) Mediation analysis on EXP2-U and EXP2-W about the effects of uncalibrated AI confidence on accuracy increase, mediated by adoption rate and disuse.
@@ 2.3.1 With Trust Calibration Support, Uncalibrated AI Lead to Distrust
Regarding trust towards AI, as depicted in Fig. 5 (e)-(g), our findings indicate that participants in EXP2-U and EXP2-O decreased their trust towards the AI’s confidence scores and overall capabilities when the AI exhibited either overconfidence or unconfidence. As the main effects of AI confidence calibration states in one-way ANOVA are significant, according to post-hoc analysis results, participants in EXP2-O had significantly lower trust towards AI’s confidence score ($p=0.004$) compared to those in EXP2-W. Similarly, participants in EXP2-U also showed significantly lower trust towards AI’s confidence score ($p=0.012$) than those in EXP2-W. Additionally, trust towards the overall capability of AI for participants in EXP2-O ($p=0.009$) was significantly lower compared to those in EXP2-W, and the same trend was observed in EXP2-U ($p=0.010$) compared to EXP2-W. These results were expected since the AI in EXP2-U and EXP2-O presented uncalibrated confidence scores.
@@ 2.3.3 Disuse Further Worsen Human-AI Collaboration Outcomes
The results of the mediation analysis for with trust calibration support group were shown in Fig. 6 (c) and (d) : Using data from EXP2-W and EXP2-O, we found that the disuse rate fully mediated the impact of AI overconfidence on participants’ adoption rate ($\beta_{i}=-10.616$, $p=0.001$; $\beta_{d}=-2.065$, $p=0.498$) and final accuracy increase ($\beta_{i}=-4.424$, $p=0.001$; $\beta_{d}=-0.101$, $p=0.926$). With data from EXP2-U and EXP2-W, the disuse rate fully mediated the effect of AI unconfidence on participants’ adoption rate ($\beta_{i}=-10.902$, $p=0.002$; $\beta_{d}=-3.056$, $p=0.291$) and final accuracy increase ($\beta_{i}=-4.357$, $p=0.002$; $\beta_{d}=-1.200$, $p=0.276$).
@@ A.2 Statistical Results
Figure 16: Linear regression results of EXP2 in Sec. 2.3.1 and Sec. 2.3.2. The symbols used to denote significance levels are as follows: * ($p<0.05$), ** ($p<0.01$), and *** ($p<0.001$).

### v3

@@ 2 Results
In both experiments, uncalibrated AI consistently hindered human-AI collaboration, and participants were unable to effectively calibrate their trust towards uncalibrated AI. Increased transparency, i.e., trust calibration support, can enhance the perception of AI’s uncalibrated state by participants and reduce misuse caused by overconfident AI; however, it introduces a distrust towards AI’s prediction accuracy as well as an increase of disuse, thereby not significantly improving the outcome of human-AI collaboration. All statistical test result details, descriptive data, additional figures, tests related to control variables, and demographic data are provided in the supplementary material.
Figure 3: Heat maps illustrating participants’ perception of AI’s confidence calibration state. Each row in a given map corresponds to different groups from EXP1 and EXP2, categorized by the confidence calibration states of the AI they collaborated with. Each column represents the AI’s confidence calibration states perceived by participants within that group, which could be unconfident, well-calibrated, or overconfident. The red boxes highlight instances where the perceived AI’s confidence calibration state matched the AI’s confidence calibration state in that group, indicating participants who correctly perceived the AI’s confidence calibration state. The numbers in each cell represent the count of participants in the experimental group who perceived the corresponding AI’s confidence calibration state for each column. Darker colors correspond to a greater number of participants. (a) is the result of EXP1. (b) is the result of EXP2.
Figure 4: One-way ANOVA results of EXP1 are depicted in violin and box plots, about the effects of AI’s confidence calibration states on the dependent variables: participants’ Adoption Rate, Disuse Rate, Misuse Rate, Accuracy Increase, Trust towards AI’s Confidence Score, Trust towards AI’s Prediction Accuracy, and Trust towards AI’s Overall Capability. The center line represents the median. The box limits represent upper and lower quartiles. The whiskers represent the 1.5x interquartile range. The points represent the outliers. The width of each violin corresponds to the frequency of observations at any given number on the y axis. The cut of kernel density estimation is set to 0. The $p$-values of the AI confidence calibration states’ main effects from the one-way ANOVA are indicated below each figure, and the significance levels from post-hoc analyses are denoted within the figures (ns: $p>0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$). F-values with degrees of freedom are also reported below the figure. Differences between groups from post-hoc analyses, excluding comparisons between EXP1-U and EXP1-O, are marked in the figure. Further information, such as the verification of homogeneity assumptions, analytical methods, and effect sizes, can be found in the supplementary materials.
@@ 2.1.1 Perceiving AI’s Overconfidence and Unconfidence is Challenging
The results of the groups EXP1-U and EXP1-O, as depicted on Fig. 3 (a), indicate that, it is difficult for the majority of participants to perceive if the provided AI’s confidence level aligns with its actual correctness likelihood. In EXP1-U, 28.571% of participants correctly perceived the AI’s unconfidence. In EXP1-O, only 26.190% of participants correctly perceived the overconfidence. The majority of participants deemed the AI’s confidence calibration state appropriate when the AI was either overconfident or unconfident, suggesting an overtrust in AI’s confidence.
@@ 2.1.4 Misuse and Disuse Further Impair Human-AI Collaboration Outcomes
The results of the mediation analysis in Fig. 6 (a) and (b) further revealed the effects of uncalibrated AI confidence on accuracy increase, mediated by adoption rate and misuse. Utilizing data from EXP1-W and EXP1-O, we found that the misuse rate ($\beta_{i}$ (indirect effect) $=-6.228$, $p<0.001$) and adoption rate ($\beta_{i}=2.854$, $p=0.007$) fully mediated the influence of AI’s overconfidence on accuracy increase ($\beta_{d}$ (direct effect) = $-1.309$, $p=0.231$). With data from EXP1-U and EXP1-W, it was discovered that the disuse rate ($\beta_{i}=-9.081$, $p<0.001$) and adoption rate ($\beta_{i}=3.196$, $p=0.003$) fully mediated the effect of AI’s unconfidence on accuracy increase ($\beta_{d}=0.487$, $p=0.612$).
Figure 5: One-way ANOVA results of EXP2 are depicted in violin and box plots, about the effects of AI’s confidence calibration states on the dependent variables: participants’ Adoption Rate, Disuse Rate, Misuse Rate, Accuracy Increase, Trust towards AI’s Confidence Score, Trust towards AI’s Prediction Accuracy, and Trust towards AI’s Overall Capability. The center line represents the median. The box limits represent upper and lower quartiles. The whiskers represent the 1.5x interquartile range. The points represent the outliers. The width of each violin corresponds to the frequency of observations at any given number on the y axis. The cut of kernel density estimation is set to 0. The $p$-values of the AI confidence calibration states’ main effects from the one-way ANOVA are indicated below each figure, and the significance levels from post-hoc analyses are denoted within the figures (ns: $p>0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$). F-values with degrees of freedom are also reported below the figure. Differences between groups from post-hoc analyses, excluding comparisons between EXP2-U and EXP2-O, are marked in the figure. Further information, such as the verification of homogeneity assumptions, analytical methods, and effect sizes, can be found in the supplementary materials.
Figure 6: The diagrams above illustrates the results of mediation analysis (ns: $p>0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$). Arrows indicate the paths of effect, while numbers represent the coefficients on each path. See more details in supplementary materials. (a) Mediation analysis on EXP1-W and EXP1-O about the effects of overconfident AI confidence on accuracy increase, mediated by adoption rate and misuse. (b) Mediation analysis on EXP1-U and EXP1-W about the effects of unconfident AI confidence on accuracy increase, mediated by adoption rate and disuse. (c) Mediation analysis on EXP2-W and EXP2-O about the effects of overconfident AI confidence on accuracy increase, mediated by adoption rate and disuse. (d) Mediation analysis on EXP2-U and EXP2-W about the effects of unconfident AI confidence on accuracy increase, mediated by adoption rate and disuse.
@@ 2.3.1 With Trust Calibration Support, Uncalibrated AI Lead to Distrust
Regarding trust towards AI, as depicted in Fig. 5 (e)-(g), our findings indicate that participants in EXP2-U and EXP2-O decreased their trust towards the AI’s confidence scores and overall capabilities when the AI exhibited either overconfidence or unconfidence. As the main effects of AI confidence calibration states in one-way ANOVA of EXP2 are significant, according to post-hoc analysis results, participants in EXP2-O had significantly lower trust towards AI’s confidence score ($p=0.004$) compared to those in EXP2-W. Similarly, participants in EXP2-U also showed significantly lower trust towards AI’s confidence score ($p=0.012$) than those in EXP2-W. Additionally, trust towards the overall capability of AI for participants in EXP2-O ($p=0.009$) was significantly lower compared to those in EXP2-W, and the same trend was observed in EXP2-U ($p=0.010$) compared to EXP2-W. These results were expected since the AI in EXP2-U and EXP2-O presented uncalibrated confidence scores.
@@ 2.3.3 Disuse Further Worsen Human-AI Collaboration Outcomes
The results of the mediation analysis in EXP2 were shown in Fig. 6 (c) and (d) : Using data from EXP2-W and EXP2-O, we found that disuse rate ($\beta_{i}=-6.969$, $p<0.001$) and adoption rate ($\beta_{i}=3.041$, $p=0.005$) fully mediated the effect of AI overconfidence on accuracy increase ($\beta_{d}=-0.596$, $p=0.459$). With data from EXP2-U and EXP2-W, disuse rate ($\beta_{i}=-6.218$, $p=0.002$) and adoption rate ($\beta_{i}=2.383$, $p=0.008$) fully mediated the effect of AI unconfidence on accuracy increase ($\beta_{d}=-1.721$, $p=0.082$).
@@ A.2 Statistical Results
Figure 16: Linear regression results of EXP2 in Sec. 2.3.1 and Sec. 2.3.2. The error bands represent a 95% confidence interval. The symbols used to denote significance levels are as follows: * ($p<0.05$), ** ($p<0.01$), and *** ($p<0.001$).